# Reward Model Overoptimisation in Iterated RLHF

### v1

@@ Acknowledgements
Appendix A Reward model comparison with the Maximum Mean Discrepancy
@@
Formally, our goal is to compare any two reward functions $R_{\phi_{1}}$ and $R_{\phi_{2}}$. As the first step, we scale both reward functions to have mean zero and variance one. This ensures that reward functions, which differ only by an affine transformation, are treated as equal to one another after scaling. For details about this result, please refer to Appendix B. This is desirable since affine transformations do not affect the ordering over policies induced by the original and transformed reward functions when they are optimised [36].
@@
As the second step, we compute the discrepancy between $R_{\phi_{1}}$ and $R_{\phi_{2}}$. While we have reward functions in principle, during training, only samples of rewards from the true and proxy are observed. Given that prompts are identically and independently distributed $x_{i}\overset{i.i.d.}{\sim}\rho$ and $y_{i}\sim\pi_{\theta}(\cdot|x_{i})$, we obtain that the observed rewards $r_{i}=R_{\phi}(x_{i},y_{i})$ are i.i.d samples (details in Appendix B). As a consequence, we can rely on the Maximum Mean Discrepancy (MMD) to measure the discrepancy between distributions of observed rewards from $R_{\phi_{1}}$ and $R_{\phi_{2}}$. The MMD compares two distributions based on their distances in the feature space determined by the chosen kernel. It is known for its strong theoretical guarantees, and it is commonly used in the two sample testing literature [14]. We use the popular squared exponential kernel.
@@ Appendix B Proofs
Let $R_{\phi_{1}},R_{\phi_{2}}\in\mathcal{R}$ be two reward functions and suppose they differ by an affine transformation, i.e. $R_{\phi_{2}}=a\cdot R_{\phi_{1}}+b$ for some $a\in\mathbb{R}^{+}$ and $b\in\mathbb{R}$. Then $R_{\phi^{\prime}_{1}}=R_{\phi^{\prime}_{2}}$, where $R_{\phi^{\prime}_{i}}=\frac{1}{\sigma_{i}}\cdot(R_{\phi_{i}}-\mu_{i})$ with $\sigma_{i}$ the standard deviation of $R_{\phi_{i}}$ and $\mu_{i}$ the mean.
@@
Given i.i.d. observations $x_{1},...,x_{n}$ from random variable $x\sim\rho$, and a policy $\pi_{\theta}$, we have that observations of rewards $r_{1},...,r_{n}$, where $r_{i}=R_{\phi}(x_{i},y_{i})$ for a deterministic reward function $R_{\phi}$ and $y_{i}\sim\pi_{\theta}(\cdot|x_{i})$ for $i=1,...,n$, are i.i.d. observations of a random variable we denote by $Z$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ C.1 Hyperparameters
Our hyperparameter settings mostly align with those used by the authors in [5]. The parameters for supervised fin-tuning are given in Table 1, reward model training hyperparameters are specified in Table 2, PPO parameters are given in Table 3, and the hyperparameters for synthesis with a policy are provided in Table 4.
@@
Learning rate $68\mathrm{e}-$ 6
@@
Learning rate $51\mathrm{e}-$ 5
@@
Learning rate $61\mathrm{e}-$ 6
Cosine annealing scheduler $71\mathrm{e}-$ 7
@@ C.2 Dataset
We use the instructions and inputs contained in the popular alpacafarm dataset [8, 38]. The entire dataset contains $52,000$ samples split into "sft" (10k), "preference" (20k), "unlabled" (20k), and "val" (2k). We use the "val" split strictly only for validation. The instructions for the reward model training are sampled from the "preference" split and the instructions for PPO are sampled from the "unlabled" split.
@@ C.3 Prompt format
We follow the prompt format used in [5, 17], which is that of the v2 format used in Open Assitant. It uses special tokens <|prompter|> and <|assistant|>, and is consistent with the GPTNeoXTokenizer class.
@@
In the case of the reward model the prompt should additionally contain an answer to the instruction, which is appended to the initial prompt and again ended with the <|endoftext|> token.
Examples for both generation and reward modelling are given in Table 5.
Table 5: Example answer generation and reward modeling prompts with proper formatting.
@@ C.4 Computational setup and cost
Appendix D Iterated RLHF and Performative Prediction
We note that the framework of performative prediction applies to our setting. In fact, when performing iterated RLHF we are simulating performative prediction or more specifically a version of strategic classification. We have that a reward model $R_{\phi}$ induces a potentially different distribution $\mathcal{D(\phi)}$ over instances $(x,y)$ where continuations $y$ are obtained from the policy $\pi_{\theta}$ optimised for $R_{\phi}$, which yields that a reward model $R_{\phi_{PO}}$ is performatively optimal if $\phi_{PO}=\underset{\phi}{\arg\min}\mathbb{E}_{(x,y)\sim\mathcal{D(\phi)}}\ell((x,y,\phi))$. Furthermore, [29] call a model $R_{\phi_{PS}}$ performatively stable if
$\phi_{PS}=\underset{\phi}{\arg\min}\mathbb{E}_{(x,y)\sim\mathcal{D}(\phi_{ps})}\ell((x,y,\phi)).$
Intuitively, retraining a performatively stable reward model after optimizing against it will yield the same reward model. As such the reward model would not be over-optimised and still perform optimally on its induced distribution. In their Theorem 3.5 the authors of [29] provide 3 conditions under which the reward model obtained from repeated iterations of RLHF converges to a unique performatively stable reward model at a linear rate. We require the loss to be $\beta$-jointly smooth and $\gamma$-strongly convex, and the map $\mathcal{D}(\cdot)$ from reward model parameters to the distribution of prompt continuation pairs to be $\epsilon$-sensitive [29]. Since as part of the map $\mathcal{D}(\cdot)$ the policy is optimised with PPO, where small changes in the reward model can lead to significant changes in the optimal policy, this mapping is generally not $\epsilon$-sensitive. As a consequence, linear convergence is not guaranteed. Note, that we may still aim for close to linear convergence by making adjustments to satisfy the stated conditions.
Appendix E Additional results
E.1 Closing the gap between proxy and gold reward function
@@
E.2 Additional results for combining preference data
@@
E.3 Additional results for reward model transfer
@@
E.4 Additional results for policy initialisation

### v2

@@ A.9 Concluding Remarks
Appendix B Reward model comparison with the Maximum Mean Discrepancy
@@
Formally, our goal is to compare any two reward functions $R_{\phi_{1}}$ and $R_{\phi_{2}}$. As the first step, we scale both reward functions to have mean zero and variance one. This ensures that reward functions, which differ only by an affine transformation, are treated as equal to one another after scaling. For details about this result, please refer to Appendix B.1. This is desirable since affine transformations do not affect the ordering over policies induced by the original and transformed reward functions when they are optimised Skalse et al. (2024).
@@
As the second step, we compute the discrepancy between $R_{\phi_{1}}$ and $R_{\phi_{2}}$. While we have reward functions in principle, during training, only samples of rewards from the true and proxy are observed. Given that prompts are identically and independently distributed $x_{i}\overset{i.i.d.}{\sim}\rho$ and $y_{i}\sim\pi_{\theta}(\cdot|x_{i})$, we obtain that the observed rewards $r_{i}=R_{\phi}(x_{i},y_{i})$ are i.i.d samples (details in Appendix B.1). As a consequence, we can rely on the Maximum Mean Discrepancy (MMD) to measure the discrepancy between distributions of observed rewards from $R_{\phi_{1}}$ and $R_{\phi_{2}}$. The MMD compares two distributions based on their distances in the feature space determined by the chosen kernel. It is known for its strong theoretical guarantees, and it is commonly used in the two sample testing literature (Gretton et al., 2012). We use the popular squared exponential kernel.
@@ B.1 Proofs
Let $R_{\phi_{1}},R_{\phi_{2}}\in\mathcal{R}$ be two reward functions and suppose they differ by an affine transformation, i.e., $R_{\phi_{2}}=a\cdot R_{\phi_{1}}+b$ for some $a\in\mathbb{R}^{+}$ and $b\in\mathbb{R}$. Then $R_{\phi^{\prime}_{1}}=R_{\phi^{\prime}_{2}}$, where $R_{\phi^{\prime}_{i}}=\frac{1}{\sigma_{i}}\cdot(R_{\phi_{i}}-\mu_{i})$ with $\sigma_{i}$ the standard deviation of $R_{\phi_{i}}$ and $\mu_{i}$ the mean.
@@
Given i.i.d. observations $x_{1},...,x_{n}$ from random variable $x\sim\rho$, and a policy $\pi_{\theta}$, we have that observations of rewards $r_{1},...,r_{n}$, where $r_{i}=R_{\phi}(x_{i},y_{i})$ for a deterministic reward function $R_{\phi}$ and $y_{i}\sim\pi_{\theta}(\cdot|x_{i})$ for i=1,,ni=1,...,n, are i.i.d. observations of a random variable we denote by $Z$.
@@
$P\left(Y_{i},Y_{j}\right)=\iint P\left(Y_{i},Y_{j}\mid X_{i},X_{j}\right)P\left(X_{i},X_{j}\right)dX_{i}dX_{j}.$
@@
$P\left(Y_{i},Y_{j}\right)=\iint P\left(Y_{i}\mid X_{i}\right)P\left(Y_{j}\mid X_{j}\right)P\left(X_{i}\right)P\left(X_{j}\right)dX_{i}dX_{j}.$
@@
$P\left(Y_{i},Y_{j}\right)=$ $\left(\int P\left(Y_{i}\mid X_{i}\right)P\left(X_{i}\right)dX_{i}\right)\times\left(\int P\left(Y_{j}\mid X_{j}\right)P\left(X_{j}\right)dX_{j}\right).$ (8)
@@
$P\left(Y_{i}=y\right)=\int P\left(Y_{i}=y\mid X_{i}=x\right)P\left(X_{i}=x\right)dx.$
@@ C.1 Hyperparameters
Our hyperparameter settings mostly align with those used by the authors in Coste et al. (2024). The parameters for supervised fin-tuning are given in Table 1, reward model training hyperparameters are specified in Table 2, PPO parameters are given in Table 3, and the hyperparameters for synthesis with a policy are provided in Table 4.
@@
Learning rate $8\mathrm{e}-6$
@@
Learning rate $1\mathrm{e}-5$
@@
Learning rate $1\mathrm{e}-6$
Cosine annealing scheduler $1\mathrm{e}-7$
@@ C.2 Dataset
We use the instructions and inputs contained in the popular AlpacaFarm dataset (Dubois et al., 2023; Taori et al., 2023). The entire dataset contains $52,000$ samples split into "sft" (10k), "preference" (20k), "unlabeled" (20k), and "val" (2k). We use the "val" split strictly only for validation. The instructions for the reward model training are sampled from the "preference" split and the instructions for PPO are sampled from the "unlabeled" split.
@@ C.3 Prompt format
We follow the prompt format used in (Coste et al., 2024; Köpf et al., 2023), which is that of the v2 format used in Open Assistant. It uses special tokens <|prompter|> and <|assistant|>, and is consistent with the GPTNeoXTokenizer class.
@@
In the case of the reward model the prompt should additionally contain an answer to the instruction, which is appended to the initial prompt and again ended with the <|endoftext|> token. Examples for both generation and reward modelling are given in Table 5.
Table 5: Example answer generation and reward modelling prompts with proper formatting.
@@ C.4 Computational setup and cost
Appendix D Additional results
D.1 Closing the gap between proxy and gold reward function
@@
D.2 Additional results for combining preference data
@@
D.3 Additional results for reward model transfer
@@
D.4 Additional results for policy initialisation
@@
D.5 On training stability across seeds and iterations
As is common with RL fine-tuning, we observed variance across random seeds. To mitigate this, we have performed training with 8 random seeds (significantly more than what is standard in the literature) and report the average performance and standard errors. While we focused on the effect of different methods on overoptimisation, we also observed that the methods proposed, particularly those that reduce overoptimisation, tend to lead to more stable training. For instance, From SFT policy initialization consistently showed lower variance in performance compared to other initialization strategies, suggesting improved stability. Please find a summary of these results in Table 6
Table 6: Mean and standard deviation across seeds at the end of the fourth iteration.
Method Mean Standard Deviation
Take last Data 0.3572 0.0406
Sample 0.2761 0.0381
Concat Data / Policy from SFT 0.4477 0.0653
Ensemble 0.3136 0.0515
Worst-Case Optimisation 0.2942 0.0450
Weight Average 0.3035 0.1248
Concat Data + LITI 0.1991 0.1678
Sample + Take last Policy -0.0632 0.1055